# On thermal transpiration and thermomolecular pressure difference

### v2

@@ 1.1. Introduction
$\lVert f\rVert_{L^{\infty}_{\alpha,0}}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ (20)
$\leq 2C_{\zeta}\max\{1,\kappa^{-1}\}\Big(\left\lVert m^{-1/2}\left(F_{in}-M\right)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\left\lVert m^{-1/2}(m_{T(x)}-m)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\Big).$
Here, $f=m^{-1/2}\left(F-M\right)$.
@@ 1.2. Organization of the article
In section 2, we review some elementary properties of the linear operator $L$. In section 3, inspired by [13], [14], and [24], we introduce a simpler and more efficient approach to obtaining the $L^{2}$ estimate and we introduce the damping term $\hat{\alpha}\nu f$ rather than $\hat{\alpha}f$ to prove Theorem 3.1 for both positive and negative $\gamma$. This approach is based on the operator $S_{\Omega}^{\kappa}K$, which is compact in $L^{2}_{0,\gamma/2}$.
@@ 2.1. Estimates for $\nu$ and $K$
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}$ $=\int_{0}^{\infty}\int_{0}^{2\pi}\int_{0}^{\pi}\frac{1}{r^{\mu-2}}e^{-\alpha_{1}r^{2}}e^{-\alpha_{3}\left(|v|\cos\theta-\alpha_{2}r\right)^{2}}\sin\theta d\theta d\phi dr$
@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
The proof can be found in [5]. We now state the main theorem of this section. The statement and strategy are similar to [13] and [14]. However, we need to modify the proof due to the difference between the mixed and diffuse boundary conditions. Moreover, we use Lemma 1.2 ([24]) to simplify the argument, and we introduce the damping term $\hat{\alpha}\nu f$ rather than $\hat{\alpha}f$ to prove Theorem 3.1 for both positive and negative $\gamma$. We present the details below.
@@
Let $\Omega\in\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Assume that $3<\gamma\leq 1$ and that (8) holds. Then, for any $\phi_{1}\in L^{2}_{0,-\gamma/2}(\Omega\times\mathbb{R}^{3})$, $\phi_{2}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{1})$, and $f_{in}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{2})$. There exists a unique solution $f\in L^{2}_{0,\gamma/2}$ to the following boundary value problem.
@@
Consider the following iteration with $f_{1}=0$,
$\left\{\begin{aligned} \hat{\alpha}\nu(v)f_{k+1}+v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})&=\phi_{1},&&\quad\text{in}\quad\in\Omega\times\mathbb{R}^{3},\\ f_{k+1}&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}&=\eta P(f_{k})+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (54)
where $0<\hat{\alpha},\eta<1$. Note that (54) and (52) are slightly different. The boundary value problem (54) includes a damping term $\hat{\alpha}\nu(v)f_{k+1}$ and we insert $\eta$ in front of $P(f)$. We first construct a solution of (54) in $L^{2}_{0,\gamma/2}$ for arbitrary $\hat{\alpha}$ and $\eta$. Let $f^{\hat{\alpha},\eta}$ denote this solution. We then show that, for fixed $\eta$, the family $\{f^{\hat{\alpha},\eta}\}$ converges in $L^{2}_{0,\gamma/2}$. Finally, we remove $\eta$ by proving that $\{f^{\eta}\}$ is a convergent sequence in $L^{2}_{0,\gamma/2}$,where $f^{\eta}=\lim_{\hat{\alpha}\to 0}f^{\hat{\alpha},\eta}$. To construct the solution of (54), by Green formula, we have
$\hat{\alpha}\lVert f_{k+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f_{k+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f_{k+1}),f_{k+1}\rangle$ (55)
@@
$\quad+C_{\hat{\alpha},\eta}\left(\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}\right)+C_{\hat{\alpha},\eta}\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
@@
$\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $+(2\hat{\alpha})^{-1}\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ (61)
@@
This implies that $\{f_{k}\}$ is a Cauchy sequence. For any fixed $\eta$ and $\hat{\alpha}$, let $\lim_{k\to\infty}f_{k}\coloneqq f^{\hat{\alpha},\eta}$. By applying Green’s formula once more, we have
$\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f^{\hat{\alpha},\eta}),f^{\hat{\alpha},\eta}\rangle$ (62)
$=\langle\phi_{1},f^{\hat{\alpha},\eta}\rangle+\frac{1}{2}\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\frac{1}{2}\lVert\eta P(f^{\hat{\alpha},\eta})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
@@
$\lVert P_{L}(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}\\ &+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.\end{pmatrix}$. (63)
@@
$\hat{\alpha}(1-aC)\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+a\lVert P_{L}(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+C_{L}\frac{1}{\kappa}\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$
@@
$+\frac{1}{2}\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\frac{1}{2}\lVert\eta P(f^{\hat{\alpha},\eta})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+\langle\phi_{1},f^{\hat{\alpha},\eta}\rangle.$
@@
$\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}$ $+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (64)
@@
Hence, $f^{\hat{\alpha},\eta}$ is a bounded sequence in $L^{2}_{0,\gamma/2}$ for any fixed $\eta$. Observe that $f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}$ satisfies
$\left\{\begin{aligned} v\cdot\nabla_{x}\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)+\frac{1}{\kappa}L\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)&=\hat{\alpha}_{1}\nu f^{\eta,\hat{\alpha}_{1}}-\hat{\alpha}_{2}\nu f^{\eta,\hat{\alpha}_{2}},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}&=\eta P\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$.
for any $0<\hat{\alpha}_{1},\hat{\alpha}_{2}<1$. After taking $\hat{\alpha}_{1}\nu f^{\eta,\hat{\alpha}_{1}}-\hat{\alpha}_{2}\nu f^{\eta,\hat{\alpha}_{2}}$ as $\phi_{1}$ and setting $f_{in}=\phi_{2}=0$, by Hölder’s inequality, we have
$\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (65)
$\leq C_{\eta}\cdot\max\{1,\kappa\}\left(\hat{\alpha}_{1}\lVert f^{\eta,\hat{\alpha}_{1}}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\hat{\alpha}_{2}\lVert f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}\right)$
@@
This implies that $f^{\eta}=\lim_{\hat{\alpha}\to 0}f^{\hat{\alpha},\eta}$ exists in $L^{2}_{0,\gamma/2}$ for any $0<\eta<1$. Moreover, applying the same procedure as in (3), we obtain
@@
$\leq C\max\{\kappa^{-2},\kappa^{2}\}\Big(\eta_{1}-\eta_{2}\Big)^{2}\lVert P(f^{\eta_{2}})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$
@@
by defining $K_{\alpha}(\cdot)\coloneq wK(w^{-1}\cdot)$. Note that $K_{\alpha}$ and $|K_{\alpha}|(f)\coloneq w\int_{\mathbb{R}^{3}}|k(v,u)|w^{-1}f(x,u)du$ have the same properties, where $k(v,u)$ denotes the kernel of $K$. In this section, we prove that problem (80) admits a unique solution in $L^{\infty}$ when $\Omega$ is unsealed with respect to $\partial\Omega_{1}$, and therefore Theorem 1.1 in subsection 4.2. First, we consider the approximating problem with the parameters $\hat{a}$ and $\eta$, and we define a recursive sequence $h^{l}$ with $h^{1}=0$.
$\left\{\begin{aligned} \hat{\alpha}\nu(v)h^{l+1}+v\cdot\nabla_{x}h^{l+1}+\frac{1}{\kappa}L_{w}(h^{l+1})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{l+1}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{l+1}&=\eta P_{w}(h^{l})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (81)
Second, we establish the following proposition to relate the $L^{\infty}$ norm of $h^{l}$ to the weighted $L^{2}$ norm of $h^{l}/w$. This relation allows us to apply the results from previous section to construct a solution to (81) in $L^{\infty}$ and then remove the parameters $\hat{a}$ and $\eta$. This approach is known as the $L^{2}-L^{\infty}$ estimate, which is an useful tool ( [13], [14]) for constructing solutions of Boltzmann equation. Finally we note that proposition 4.1 remains valid when $\hat{a}=0$ and $\eta=1$. The statement of the proposition is
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\frac{C_{\zeta}}{\kappa}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}w^{-1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$

### v3

@@ Abstract
Keyword : Stationary Boltzmann equation, thermal transpiration, thermomocular pressure difference, mixed boundary condition.
@@ 1.1. Introduction
$\lVert m^{-1/2}\left(F-M\right)\rVert_{L^{\infty}_{\alpha,0}}+\lVert m^{-1/2}\left(F-M\right)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ (20)
$\leq 2C_{\zeta}\Big(\left\lVert m^{-1/2}\left(F_{in}-M\right)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\left\lVert m^{-1/2}(m_{T(x)}-m)\right\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\Big).$
@@ 1.2. Organization of the article
In section 2, we review some elementary properties of the linear operator $L$. In section 3, inspired by [14], [15], and [25], we introduce a simpler and more efficient approach to obtaining the $L^{2}$ estimate This approach is based on the operator $S_{\Omega}^{\kappa}K$, which is compact in $L^{2}_{0,\gamma/2}$.
@@ 2.1. Estimates for $\nu$ and $K$
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}du$ $=\int_{0}^{\infty}\int_{0}^{2\pi}\int_{0}^{\pi}\frac{1}{r^{\mu-2}}e^{-\alpha_{1}r^{2}}e^{-\alpha_{3}\left(|v|\cos\theta-\alpha_{2}r\right)^{2}}\sin\theta d\theta d\phi dr$
@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
The proof can be found in [6]. We now state the main theorem of this section. The statement and strategy are similar to [14] and [15]. However, we need to modify the proof due to the difference between the mixed and diffuse boundary conditions. Moreover, we use Lemma 1.2 ([25]) to simplify the argument. We present the details below.
@@
Let $\Omega\in\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Assume that $3<\gamma\leq 1$ and that (8) holds. Then, for any $\phi_{1}\in L^{2}_{0,\max\{0,-\gamma/2\}}(\Omega\times\mathbb{R}^{3})$, $\phi_{2}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{1})$, and $f_{in}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{2})$. There exists a unique solution $f\in L^{2}_{0,\gamma/2}$ to the following boundary value problem.
@@
We consider the following iteration with $f_{1}=0$,
$\left\{\begin{aligned} \varpi f_{k+1}+v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})&=\phi_{1},&&\quad\text{in}\quad\in\Omega\times\mathbb{R}^{3},\\ f_{k+1}&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}&=\eta P(f_{k})+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (54)
where $0<\varpi,\eta<1$. Note that (54) differs slightly (52). More precisely, the boundary value problem (54) contains a penalization term $\varpi f_{k+1}$ and the factor $\eta$ inserted in front of $P(f)$. We first construct a solution to (54) in $L^{2}_{0,\gamma/2}$ for arbitrary $\varpi$ and $\eta$ and denote this solution by $f^{\varpi,\eta}$. We then show that, for fixed $\eta$, the family $\{f^{\varpi,\eta}\}$ converges in $L^{2}_{0,\gamma/2}$. Finally, we remove $\eta$ by proving that $\{f^{\eta}\}$ is a convergent sequence in $L^{2}_{0,\gamma/2}$,where $f^{\eta}=\lim_{\varpi\to 0}f^{\varpi,\eta}$. We emphasize that the penalization term is introduced in order to justify the application of Green’s formula to (54) for all $3<\gamma\leq 1$. Consequently, when $\gamma<0$, the procedure for removing $\varpi$ and $\eta$ are different from that in the case of $\gamma\geq 0$. Since the argument for the negative $\gamma$ case relies on an inequality established in section 4, for the sake of clarity we present only the case $\gamma\geq 0$ here. The case $\gamma<0$ can be treated by applying the same argument as in [14]. To construct the solution of (54), by Green formula, we have
$\varpi\lVert f_{k+1}\rVert^{2}_{L^{2}}+\frac{1}{2}\lVert f_{k+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f_{k+1}),f_{k+1}\rangle$ (55)
@@
$\quad+C_{\varpi,\eta}\left(\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}\right)+C_{\varpi,\eta}\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
@@
$\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}}$ $+(2\varpi)^{-1}\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ (61)
@@
This implies that $\{f_{k}\}$ is a Cauchy sequence. For any fixed $\eta$ and $\varpi$, let $\lim_{k\to\infty}f_{k}\coloneqq f^{\varpi,\eta}$. By applying Green’s formula once more, we have
$\varpi\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}}+\frac{1}{2}\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f^{\varpi,\eta}),f^{\varpi,\eta}\rangle$ (62)
$=\langle\phi_{1},f^{\varpi,\eta}\rangle+\frac{1}{2}\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\frac{1}{2}\lVert\eta P(f^{\varpi,\eta})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
@@
$\lVert P_{L}(f^{\varpi,\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\varpi\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}}+\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f^{\varpi,\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}\\ &+\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\varpi,\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.\end{pmatrix}$. (63)
@@
$\varpi(1-aC)\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}}+a\lVert P_{L}(f^{\varpi,\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+C_{L}\frac{1}{\kappa}\lVert(I-P_{L})(f^{\varpi,\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$
@@
$+\frac{1}{2}\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\frac{1}{2}\lVert\eta P(f^{\varpi,\eta})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+\langle\phi_{1},f^{\varpi,\eta}\rangle.$
@@
$\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\varpi,\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}$ $+\lVert(I-P)f^{\varpi,\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f^{\varpi,\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (64)
@@
Hence, $f^{\varpi,\eta}$ is a bounded sequence in $L^{2}_{0,\gamma/2}$ for any fixed $\eta$. Observe that $f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}$ satisfies
$\left\{\begin{aligned} v\cdot\nabla_{x}\Big(f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}\Big)+\frac{1}{\kappa}L\Big(f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}\Big)&=\varpi_{1}f^{\eta,\varpi_{1}}-\varpi_{2}f^{\eta,\varpi_{2}},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}&=\eta P\Big(f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}\Big),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$.
for any $0<\varpi_{1},\varpi_{2}<1$. After taking $\varpi_{1}f^{\eta,\varpi_{1}}-\varpi_{2}f^{\eta,\varpi_{2}}$ as $\phi_{1}$ and setting $f_{in}=\phi_{2}=0$, by Hölder’s inequality, we have
$\lVert f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta,\varpi_{1}}-f^{\eta,\varpi_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (65)
$\leq C_{\eta}\cdot\kappa\left(\varpi_{1}\lVert f^{\eta,\varpi_{1}}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\varpi_{2}\lVert f^{\eta,\varpi_{2}}\rVert^{2}_{L^{2}_{0,-\gamma/2}}\right)$
@@
This implies that $f^{\eta}=\lim_{\varpi\to 0}f^{\varpi,\eta}$ exists in $L^{2}_{0,\gamma/2}$ for any $0<\eta<1$. Moreover, applying the same procedure as in (3), we obtain
@@
$\leq C\max\{1,\kappa^{2}\}\Big(\eta_{1}-\eta_{2}\Big)^{2}\lVert P(f^{\eta_{2}})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$
@@
by defining $K_{\alpha}(\cdot)\coloneq wK(w^{-1}\cdot)$. Note that $K_{\alpha}$ and $|K_{\alpha}|(f)\coloneq w\int_{\mathbb{R}^{3}}|k(v,u)|w^{-1}f(x,u)du$ have the same properties, where $k(v,u)$ denotes the kernel of $K$. In this section, we prove that problem (80) admits a unique solution in $L^{\infty}$ when $\Omega$ is unsealed with respect to $\partial\Omega_{1}$, and therefore Theorem 1.1 in subsection 4.2. First, we consider the approximating problem with the parameters $\varpi$ and $\eta$, and we define a recursive sequence $h^{l}$ with $h^{1}=0$.
$\left\{\begin{aligned} \varpi h^{l+1}+v\cdot\nabla_{x}h^{l+1}+\frac{1}{\kappa}L_{w}(h^{l+1})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{l+1}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{l+1}&=\eta P_{w}(h^{l})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (81)
Second, we establish the following proposition to relate the $L^{\infty}$ norm of $h^{l}$ to the weighted $L^{2}$ norm of $h^{l}/w$. This relation allows us to apply the results from previous section to construct a solution to (81) in $L^{\infty}$ and then remove the parameters $\varpi$ and $\eta$. This approach is known as the $L^{2}-L^{\infty}$ estimate, which is an useful tool ( [14], [15]) for constructing solutions of Boltzmann equation. Finally we note that proposition 4.1 remains valid when $\varpi=0$ and $\eta=1$. The statement of the proposition is
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\min\{1,\kappa^{-1}\}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}w^{-1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$